OpenStudy (anonymous):

4x+2x+4=5x-5 pleasee help...

OpenStudy (anonymous):

6x+4 = 5x-5 -5x -5x 1x+4= -5 -4 -4 x=-9

OpenStudy (leelee):

x=-9